OpenStudy (anonymous):

5u+3v= -15 2u+ v= -5

OpenStudy (anonymous):

Equation: 5u+3v = -15 a = 5 b = 3 c = -15 Subtracting 3 from both sides, we get: 5u = -18 -3 /5 or u = -3.6 Equation: 2x+0 = -5 a = 2 b = 0 c = -5 Subtracting 0 from both sides, we get: 2u = -5 0 /2 or u = -2.5